Consider an economy where the marginal propensity to consume (MPC) for the average person is 0.5. If the government is in deep debt (think about the situation in the U.S) and decides to increase the tax by 1 billion dollars to pay off some of its debt. How would the GDP respond to the tax increase in the Goods Market Equilibrium? Expain.
Answer #1

Goods market equilibrium shows Y = C+I+G+NX, where C = Co + CYd , where Yd = Disposable income, Due to tax increase the multiplier will be - T/ 1- MPC, here MPC= 0.5, Tax = 1 billion. GDP will change by = -1*2 = -2 billion. As 1/1- 0.5 = 1/0.5 =2. This tax increase will reduce the GDP by tax increase multiplied by multiplier and it will be 2 billion. So GDP will decrease by 2 billion.